OpenStudy (anonymous):

Solve. 2x^2 + x = 15.

OpenStudy (waleed_imtiaz):

U can write it as 2x^2+x-15=0 2x^2+6x-5x-15=0 (pairing) take 2x as common from the first two terms and take 5 as common from the last two terms... 2x(x+3)-5(x+3)=0 (x+3)(2x-5)=0 so x=-3 and x=5/2

OpenStudy (anonymous):

oh okay ! thanks .

OpenStudy (waleed_imtiaz):

U're welcome.....